# Score Change of Variables

### v1

@@ 2.2 Score Change of Variables
$\nabla_{\mathbf{y}}\log q(\mathbf{y})=\mathbf{J}_{\phi^{-1}}(\mathbf{y})^{\top}\nabla_{\mathbf{x}}\log p(\mathbf{x})+\nabla_{\mathbf{x}}\cdot\left(\mathbf{J}_{\phi^{-1}}(\phi(\mathbf{x}))^{\top}\right)\bigg{|}_{\mathbf{x}=\phi^{-1}(\mathbf{y})}.$
@@ 2.3 Special Cases
$\nabla_{y}\log q(y)=\frac{1}{(\phi^{\prime}(x))^{2}}\left(\phi^{\prime}(x)\nabla_{x}\log p(x)-\phi^{\prime\prime}(x)\right)\bigg{|}_{x=\phi^{-1}(y)}.$
@@
$\nabla_{y}\log q(y)=\frac{1}{\|\nabla v(\mathbf{x})\|_{2}^{2}}\left(\nabla v(\mathbf{x})^{\top}\nabla_{\mathbf{x}}\log p(\mathbf{x})-\Delta v(\mathbf{x})-\sum_{i=1}^{n}\left(\frac{\partial v}{\partial x_{i}}\right)^{2}\frac{\partial}{\partial y}\log p((x_{j})_{j\neq i}\mid y)\right).$
@@ 3.1 Reverse-Time Itô Lemma
where $\mathbf{W}(t)$ is a standard Wiener process, the corresponding reverse-time SDE [1, 25] is:
$\mathrm{d}\mathbf{X}$ $=\left[\mathbf{f}(\mathbf{X},t)-\mathbf{G}(\mathbf{X},t)\mathbf{G}(\mathbf{X},t)^{\top}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{X})-\nabla_{\mathbf{x}}\cdot\left(\mathbf{G}(\mathbf{X},t)\mathbf{G}(\mathbf{X},t)^{\top}\right)\right]\mathrm{d}t$ (3)
$\quad+\mathbf{G}(\mathbf{X},t)\,\mathrm{d}\bar{\mathbf{W}}(t),$ (4)
where $\bar{\mathbf{W}}(t)$ is a Wiener process in reverse time. This reverse process allows us to recover the original data distribution by starting from noise.
In many scenarios, we prefer to perform this generative diffusion process in a transformed space, for instance when data lies on a manifold [18, 6]. Consider a bijective, twice continuously differentiable transformation $\phi:\mathbb{R}^{n}\times[0,T]\to\mathbb{R}^{n}$. The transformed process $\mathbf{Y}(t)=\phi(\mathbf{X}(t),t)$ for $t\in[0,T]$ satisfies a new SDE [17]:
$\mathrm{d}\mathbf{Y}=\bar{\mathbf{f}}(\mathbf{Y},t)\,\mathrm{d}t+\bar{\mathbf{G}}(\mathbf{Y},t)\,\mathrm{d}\mathbf{W}(t),$ (5)
where the transformed drift $\bar{\mathbf{f}}$ and diffusion $\bar{\mathbf{G}}$ are obtained via Itô’s Lemma:
$\bar{\mathbf{f}}(\mathbf{y},t)$ $=\left(\frac{\partial\phi(\mathbf{x},t)}{\partial t}+\mathbf{J}_{\phi}(\mathbf{x},t)\mathbf{f}(\mathbf{x},t)+\frac{1}{2}\text{Tr}\left[\mathbf{G}(\mathbf{x},t)^{\top}\mathbf{H}_{\phi}(\mathbf{x},t)\mathbf{G}(\mathbf{x},t)\right]\right)\bigg{|}_{\mathbf{x}=\phi^{-1}(\mathbf{y},t)},$
$\bar{\mathbf{G}}(\mathbf{y},t)$ $=\mathbf{J}_{\phi}(\mathbf{x},t)\mathbf{G}(\mathbf{x},t)\bigg{|}_{\mathbf{x}=\phi^{-1}(\mathbf{y},t)},$
and $\mathbf{H}_{\phi}$ is the Hessian tensor of $\phi$ with respect to $\mathbf{x}$.
@@
Let $\phi:\mathbb{R}^{n}\times[0,\infty)\to\mathbb{R}^{n}$ be a bijective and twice-differentiable transformation for each $t$. Let $p_{t}(\mathbf{x})$ and $q_{t}(\mathbf{y})$ be the probability density functions at time $t$ of $\mathbf{X}$ and $\mathbf{Y}=\phi(\mathbf{X},t)$, respectively. The reverse-time SDE of $\mathbf{Y}$ is:
$\mathrm{d}\mathbf{Y}$ $=\left[\bar{\mathbf{f}}(\mathbf{Y},t)-\nabla_{\mathbf{y}}\cdot\bigl{(}\bar{\mathbf{G}}(\mathbf{Y},t)\bar{\mathbf{G}}(\mathbf{Y},t)^{\top}\bigr{)}\right.$
$\quad\left.-\bar{\mathbf{G}}(\mathbf{Y},t)\bar{\mathbf{G}}(\mathbf{Y},t)^{\top}\left(\mathbf{J}_{\phi^{-1}}(\mathbf{Y},t)^{\top}\nabla_{\mathbf{x}}\log p_{t}(\phi^{-1}(\mathbf{Y},t))\right.\right.$
$\quad\quad\left.\left.+\nabla_{\mathbf{x}}\cdot\left(\mathbf{J}_{\phi^{-1}}(\phi(\mathbf{X},t),t)^{\top}\right)\big{|}_{\mathbf{x}=\phi^{-1}(\mathbf{y},t)}\right)\right]\mathrm{d}t+\bar{\mathbf{G}}(\mathbf{Y},t)\mathrm{d}\bar{\mathbf{W}}(t).$ (6)
This result enables sampling in the transformed space while using score functions learned in the original space. In practice, we approximate $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ with a neural network $s_{\theta}(\mathbf{x},t)$ trained via denoising score matching [26]. To account for the transformation, we modify the training loss to incorporate appropriate Jacobian weights (see Appendix A.4 for details):
$L(\theta)=\mathbb{E}_{t}\left[\lambda(t)\,\mathbb{E}_{\mathbf{X}(0)}\,\mathbb{E}_{\mathbf{X}|\mathbf{X}(0)}\left\|\mathbf{J}_{\phi^{-1}}(\phi(\mathbf{X},t),t)^{\top}\left(s_{\theta}(\mathbf{X},t)-\nabla_{\mathbf{x}}\log p_{0t}(\mathbf{X}|\mathbf{X}(0))\right)\right\|^{2}\right]$ (7)
@@ 3.2.3 Generalized Sliced Score Matching
$\quad+\mathbb{E}_{p_{d}}\mathbb{E}_{p_{v}}\left[s_{\theta}(\mathbf{X})^{\top}\mathbf{H}_{v}(\mathbf{X})\nabla_{\mathbf{x}}v(\mathbf{X})\right]+\mathbb{E}_{p_{d}}\mathbb{E}_{p_{v}}\left[\nabla_{\mathbf{x}}v(\mathbf{X})^{\top}s_{\theta}(\mathbf{X})\Delta v(\mathbf{X})\right].$ (11)
@@ 3.2.4 Derivation Outline
To derive GSSM, we first apply standard score matching in the transformed space $y=v(\mathbf{x})$. Then, using Corollary 2, we substitute the transformed score back into the original space. Integration by parts under appropriate conditions (see Appendix B) yields Equation 3.2.3. The complete derivation is in Appendix B.
@@ 4.1 Diffusion on the Probability Simplex for Chess Positions
We demonstrate our reverse-time Itô lemma by generating chess positions through diffusion on a probability simplex. Prior work has explored diffusion on the simplex by training directly in a constrained space [12], but our approach enables training in an unconstrained space with a Gaussian score model while still sampling in the simplex.
@@ 4.1.2 Transformation Between Spaces
$\psi_{i}(\mathbf{p})$ $=\log\left(\frac{p_{i}}{1-\sum_{j=1}^{12}p_{j}}\right).$ (13)
@@ 4.1.3 Training and Sampling
We train a score-based model $s_{\theta}(\mathbf{x},t)$ in $\mathbb{R}^{12}$. Chess positions are represented with a softened one-hot encoding to avoid infinite values when applying the inverse transformation $\psi(\cdot)$. The score model is trained using a Variance Preserving (VP) SDE [25]:
@@
where $\beta(t)$ is the noise schedule. During inference, we apply our reverse-time Itô lemma to perform diffusion directly in the projected simplex. By leveraging the learned scores in the unconstrained space and the known transformation $\phi$, we obtain the appropriate transformed SDE coefficients (derived in Appendix E) and sample valid chess positions on the simplex.
@@ 4.1.4 Controlling Piece Density
The transformed drift term, $\bar{\mathbf{f}}(\mathbf{Y},t)$, interacts with the simplex geometry to influence piece distributions. To gain intuitive control, we introduce a scaling factor $w$:
$\tilde{\mathbf{f}}(\mathbf{Y},t)=w\bar{\mathbf{f}}(\mathbf{Y},t).$ (15)
Increasing $w$ biases the distribution toward fewer pieces (increasing empty squares), while decreasing it encourages more pieces. This simple scalar parameter provides interpretable geometric control over the final configurations, allowing practitioners to shape the sampling process easily. The generated positions in Figure 5 demonstrate how varying $w$ changes piece density without losing consistency in the underlying chess position representation.
Figure 1: Generated positions with $w=0.9$.
Figure 2: Generated positions with $w=1.0$.
Figure 3: Generated positions with $w=1.1$.
Figure 4: Generated positions with $w=1.2$.
Figure 5: Generated chess positions with different values of $w$. Increasing $w$ reduces the number of pieces on the board, illustrating the intuitive geometric control enabled by sampling in the transformed simplex space.
@@ 4.2.2 Variance Reduction
$\mathcal{L}_{\text{GSSM-VR}}(s_{\theta})$ $=\frac{1}{2}\mathbb{E}_{p_{d}}\Big{[}\left(\sigma_{1}^{2}-2\sigma_{2}^{2}\right)\sum_{i}s_{\theta i}(\mathbf{X})^{2}x_{i}^{2}+\sigma_{2}^{2}\left(\|\mathbf{X}\|^{2}\|s_{\theta}(\mathbf{X})\|^{2}+(s_{\theta}(\mathbf{X})^{\top}\mathbf{X})^{2}\right)+\sigma_{3}^{2}\|s_{\theta}(\mathbf{X})\|^{2}\Big{]}$ (19)
$\quad+\mathbb{E}_{p_{d}}\left[\left(2\sigma_{1}^{2}+(n-1)\sigma_{2}^{2}\right)s_{\theta}(\mathbf{X})^{\top}\mathbf{X}\right]$
$\quad+\mathbb{E}_{p_{d}}\mathbb{E}_{\mathbf{A},\mathbf{b}}\left[(\mathbf{A}\mathbf{X}+\mathbf{b})^{\top}\nabla_{\mathbf{x}}s_{\theta}(\mathbf{X})(\mathbf{A}\mathbf{X}+\mathbf{b})\right].$
The full derivation for the quadratic case is provided in Appendix C.
@@ 4.2.3 Score Function Representation in DKEF Models
Additional details on the DKEF model architecture are provided in Appendix D.
@@ 4.2.4 Experimental Setup
We evaluate SM, SSM, GSSM, and their variance-reduced variants on three UCI datasets (Parkinsons, RedWine, WhiteWine) [2], following the training protocol and hyperparameters from [23]. For the quadratic transformation, we set $\sigma_{3}^{2}=1$ and choose the matrix variances as $\sigma_{2}^{2}=1/\sqrt{n}$ and $\sigma_{1}^{2}=2/\sqrt{n}$, where $n$ is the input dimension. The vector $\mathbf{b}$ is sampled from a Rademacher distribution, and entries of $\mathbf{A}$ are sampled from a Gaussian distribution with the given variances. This setup allows efficient sampling of $\mathbf{A}\mathbf{x}$ without explicitly constructing $\mathbf{A}$ (see Appendix C). For SSM-VR and SSM, we sample the random projection vectors $\mathbf{v}$ from a Rademacher distribution, following [23].

### v2

@@ 2.2 Score Change of Variables
$\nabla_{\mathbf{y}}\log q(\mathbf{y})=\mathbf{J}_{\phi^{-1}}(\mathbf{y})^{\top}\nabla_{\mathbf{x}}\log p(\mathbf{x})\nabla_{\mathbf{x}}\cdot\left(\mathbf{J}_{\phi^{-1}}(\phi(\mathbf{x}))^{\top}\right)\bigg{|}_{\mathbf{x}=\phi^{-1}(\mathbf{y})}.$
@@
  • We will frequently work with transformations and their inverses. When a transformation $\mathbf{y}=\phi(\mathbf{x})$ is defined, we may drop the explicit evaluation notation $\big{|}_{\mathbf{x}=\phi^{-1}(\mathbf{y})}$, and it will be implicitly understood that $\mathbf{x}$ is substituted with $\phi^{-1}(\mathbf{y})$ within subsequent formulas, unless otherwise specified.
@@ 2.3 Special Cases
$\nabla_{y}\log q(y)=\frac{1}{(\phi^{\prime}(x))^{2}}\left(\phi^{\prime}(x)\nabla_{x}\log p(x)-\phi^{\prime\prime}(x)\right).$
@@
$\nabla_{y}\log q(y)={}$ $\frac{1}{\|\nabla v(\mathbf{x})\|_{2}^{2}}\Big{(}\nabla v(\mathbf{x})^{\top}\nabla_{\mathbf{x}}\log p(\mathbf{x})-\Delta v(\mathbf{x})-\sum_{i=1}^{n}\left(\frac{\partial v}{\partial x_{i}}\right)^{2}\frac{\partial}{\partial y}\log p((x_{j})_{j\neq i}\mid y)\Big{)}.$
@@ 3.1 Reverse-Time Itô Lemma
where $\mathbf{W}(t)$ is a standard Wiener process, we define the reverse drift:
$\bar{\mathbf{f}}(\mathbf{x},t)={}$ $\mathbf{f}(\mathbf{x},t)-\mathbf{G}(\mathbf{x},t)\mathbf{G}(\mathbf{x},t)^{\top}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})-\nabla_{\mathbf{x}}\cdot\left(\mathbf{G}(\mathbf{x},t)\mathbf{G}(\mathbf{x},t)^{\top}\right).$ (3)
The corresponding reverse-time SDE [1, 31] evolves backward in time from $t=T$ to $t=0$. Let $\bar{\mathbf{X}}_{t}$ denote the reverse process state at time $t$, where $t$ decreases from $T$ to 00. The differential $\mathrm{d}t$ now represents an infinitesimal negative time increment:
$\mathrm{d}\bar{\mathbf{X}}$ $=\bar{\mathbf{f}}(\bar{\mathbf{X}},t)\,\mathrm{d}t+\mathbf{G}(\bar{\mathbf{X}},t)\,\mathrm{d}\bar{\mathbf{W}}(t),$ (4)
where $\bar{\mathbf{W}}(t)$ is a reverse-time Wiener process (adapted to a decreasing filtration).
In many scenarios, we prefer to perform this generative diffusion process in a transformed space or with a more complex forward process, for instance when data lies on a manifold [22, 7]. Let $\phi:\mathbb{R}^{n}\times[0,T]\to\mathbb{R}^{n}$ be a bijective $C^{2,1}$ transformation. The transformed process $\mathbf{Y}(t)=\phi(\mathbf{X}(t),t)$ for $t\in[0,T]$ satisfies a new SDE [20]:
$\mathrm{d}\mathbf{Y}=\tilde{\mathbf{f}}(\mathbf{Y},t)\,\mathrm{d}t+\tilde{\mathbf{G}}(\mathbf{Y},t)\,\mathrm{d}\mathbf{W}(t),$ (5)
where the transformed drift $\tilde{\mathbf{f}}$ and diffusion $\bar{\mathbf{G}}$ are obtained via Itô’s Lemma. For $\mathbf{x}=\phi^{-1}(\mathbf{y},t)$:
$\tilde{\mathbf{f}}(\mathbf{y},t)={}$ $\frac{\partial\phi(\mathbf{x},t)}{\partial t}+\mathbf{J}_{\phi}(\mathbf{x},t)\mathbf{f}(\mathbf{x},t)+\frac{1}{2}\text{Tr}\left[\mathbf{G}(\mathbf{x},t)^{\top}\mathbf{H}_{\phi}(\mathbf{x},t)\mathbf{G}(\mathbf{x},t)\right],$ (6)
$\tilde{\mathbf{G}}(\mathbf{y},t)={}$ $\mathbf{J}_{\phi}(\mathbf{x},t)\mathbf{G}(\mathbf{x},t),$ (7)
where $\mathbf{H}_{\phi}$ is the Hessian tensor of $\phi$ with respect to $\mathbf{x}$.
@@
Let $\phi:\mathbb{R}^{n}\times[0,T]\to\mathbb{R}^{n}$ be a bijective $C^{2,1}$ transformation. Let $p_{t}(\mathbf{x})$ and $q_{t}(\mathbf{y})$ be the probability density functions at time $t$ of $\mathbf{X}$ and $\mathbf{Y}=\phi(\mathbf{X},t)$, respectively. The reverse-time SDE of $\mathbf{Y}$ is:
$\mathrm{d}\bar{\mathbf{Y}}=\hat{\mathbf{f}}(\bar{\mathbf{Y}},t)dt+\tilde{\mathbf{G}}(\mathbf{\bar{Y}},t)\mathrm{d}\bar{\mathbf{W}}(t).$ (8)
where the diffusion $\tilde{\mathbf{G}}$ is defined in (7) and the reverse drift $\hat{\mathbf{f}}$ is given by:
$\hat{\mathbf{f}}(\mathbf{y},t)$ $=\frac{\partial\phi(\mathbf{x},t)}{\partial t}+\mathbf{J}_{\phi}(\mathbf{x},t)\bar{\mathbf{f}}(\mathbf{x},t)-\frac{1}{2}\text{Tr}\left[\mathbf{G}(\mathbf{x},t)^{\top}\mathbf{H}_{\phi}(\mathbf{x},t)\mathbf{G}(\mathbf{x},t)\right].$ (9)
This result can be interpreted through two complementary perspectives:
  1. 1.
    Reversing the Transformed Process: Directly applying the time-reversal formula (3) to the forward SDE of $\mathbf{Y}$ in (5), substituting $\tilde{\mathbf{f}}$ and $\tilde{\mathbf{G}}$ into (3).
  2. 2.
    Transforming the Reversed Process: Applying Itô’s Lemma to the original reverse process $\bar{\mathbf{X}}_{t}$ (4) under $\phi(\cdot,t)$.
We show in Appendix B that both approaches simplify to (8) for $\bar{\mathbf{Y}}_{t}$, confirming the path-wise equivalence:
$\text{Reverse}(\phi(\mathbf{X}_{t},t))\equiv\phi(\text{Reverse}(\mathbf{X}_{t}),t).$
In practice, we approximate $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ with a neural network $s_{\theta}(\mathbf{x},t)$ trained via denoising score matching [32]. To account for the transformation, we modify the training loss to incorporate appropriate Jacobian weights (see Appendix B.5 for details):
$L(\theta)$ $=\mathbb{E}_{t}\Big{[}\lambda(t)\,\mathbb{E}_{\mathbf{X}(0)}\,\mathbb{E}_{\mathbf{X}|\mathbf{X}(0)}\Big{\|}\mathbf{J}_{\phi^{-1}}(\phi(\mathbf{X},t),t)^{\top}\Big{(}s_{\theta}(\mathbf{X},t)-\nabla_{\mathbf{x}}\log p_{0t}(\mathbf{X}|\mathbf{X}(0))\Big{)}\Big{\|}^{2}\Big{]}$ (10)
@@ 3.2.3 Generalized Sliced Score Matching
$\quad+\mathbb{E}_{p_{d}}\mathbb{E}_{p_{v}}\left[s_{\theta}(\mathbf{X})^{\top}\mathbf{H}_{v}(\mathbf{X})\nabla_{\mathbf{x}}v(\mathbf{X})\right]\mathbb{E}_{p_{d}}\mathbb{E}_{p_{v}}\left[\nabla_{\mathbf{x}}v(\mathbf{X})^{\top}s_{\theta}(\mathbf{X})\Delta v(\mathbf{X})\right].$ (14)
@@ 3.2.4 Derivation Outline
To derive GSSM, we first apply standard score matching in the transformed space $y=v(\mathbf{x})$. Then, using Corollary 2, we substitute the transformed score back into the original space. Integration by parts under appropriate conditions (see Appendix C) yields Equation 3.2.3. The complete derivation is in Appendix C.
@@ 4.1 Diffusion on the Probability Simplex for Chess Positions
We demonstrate our reverse-time Itô lemma by generating chess positions through diffusion on a probability simplex. Prior work has explored the same diffusion process on the simplex by training directly in a constrained space [14], but our approach enables training in an unconstrained space with a Gaussian score model while still sampling in the simplex.
@@ 4.1.2 Transformation Between Spaces
$\phi^{-1}_{i}(\mathbf{y})$ $=\log\left(\frac{y_{i}}{1-\sum_{j=1}^{12}y_{j}}\right).$ (16)
@@ 4.1.3 Training and Sampling
We train a score-based model $s_{\theta}(\mathbf{x},t)$ in $\mathbb{R}^{12}$. Chess positions are represented with a softened one-hot encoding to avoid infinite values when applying the inverse transformation. The score model is trained using a Variance Preserving (VP) SDE [31]:
@@
where $\beta(t)$ is the noise schedule. During inference, we apply the reverse-time Itô lemma to perform diffusion directly in the projected simplex. By leveraging the learned scores in the unconstrained space and the known transformation $\phi$, we obtain the appropriate transformed SDE coefficients (derived in Appendix F) and sample valid chess positions on the simplex.
@@ 4.1.4 Controlling Piece Density
The transformed reverse drift term, $\hat{\mathbf{f}}(\mathbf{Y},t)$, interacts with the simplex geometry to influence piece distributions. To gain intuitive control, we introduce a scaling factor $w$:
$\hat{\mathbf{f}}_{w}(\mathbf{Y},t)=w\hat{\mathbf{f}}(\mathbf{Y},t).$ (18)
Decreasing $w$ biases the distribution toward fewer pieces (increasing empty squares), while increasing it encourages more pieces. This scalar parameter provides interpretable geometric control over the final configurations, allowing practitioners to shape the sampling process easily. The generated positions in Figure 5 demonstrate how varying $w$ changes piece density without losing consistency in the underlying chess position representation.
Figure 1: $w=0.8$.
Figure 2: $w=0.9$.
Figure 3: $w=1.0$.
Figure 4: $w=1.1$.
Figure 5: Generated chess positions with different values of $w$. Increasing $w$ increases the number of pieces on the board.
@@ 4.2.2 Variance Reduction
$\mathcal{L}_{\text{GSSM-VR}}(s_{\theta})=\mathbb{E}_{p_{d}}\Big{[}$ $\frac{1}{2}L_{1}+L_{2}+\mathbb{E}_{\mathbf{A},\mathbf{b}}\left[L_{3}\right]\Big{]},$ (22)
where
$L_{1}$ $=(\sigma_{1}^{2}-2\sigma_{2}^{2})\sum_{i}s_{\theta i}(\mathbf{X})^{2}x_{i}^{2}+\sigma_{2}^{2}(\|\mathbf{X}\|^{2}\|s_{\theta}(\mathbf{X})\|^{2}+(s_{\theta}(\mathbf{X})^{\top}\mathbf{X})^{2})+\sigma_{3}^{2}\|s_{\theta}(\mathbf{X})\|^{2}$
$L_{2}$ $=(2\sigma_{1}^{2}+(n-1)\sigma_{2}^{2})s_{\theta}(\mathbf{X})^{\top}\mathbf{X}$
$L_{3}$ $=(\mathbf{A}\mathbf{X}+\mathbf{b})^{\top}\nabla_{\mathbf{x}}s_{\theta}(\mathbf{X})(\mathbf{A}\mathbf{X}+\mathbf{b})$
The full derivation for the quadratic case is provided in Appendix D.
@@ 4.2.3 Score Function Representation in DKEF Models
Additional details on the DKEF model architecture are provided in Appendix E.
@@ 4.2.4 Experimental Setup
We evaluate SM, SSM, GSSM, and their variance-reduced variants on three UCI datasets (Parkinsons, RedWine, WhiteWine) [2], following the training protocol and hyperparameters from [29]. For the quadratic transformation, we set $\sigma_{3}^{2}=1$ and choose the matrix variances as $\sigma_{2}^{2}=1/\sqrt{n}$ and $\sigma_{1}^{2}=2/\sqrt{n}$, where $n$ is the input dimension. The vector $\mathbf{b}$ is sampled from a Rademacher distribution, and entries of $\mathbf{A}$ are sampled from a Gaussian distribution with the given variances. This setup allows efficient sampling of $\mathbf{A}\mathbf{x}$ without explicitly constructing $\mathbf{A}$ (see Appendix D). For SSM-VR and SSM, we sample the random projection vectors $\mathbf{v}$ from a Rademacher distribution, following [29].